A square has an area of 225 square feet. Find the length of each side.
Answer by jim_thompson5910(35256)

A = s^2


225 = s^2


s^2 = 225


s = sqrt(225)


s = 15


The side length is 15 feet
------------------------------------------------------------------------

Let me know if that helps or not. Thanks.

If you need more help, feel free to email me at jim_thompson5910@hotmail.com

My Website: http://www.freewebs.com/jimthompson5910/home.html